# sample.bib

### v1

@@ Abstract
Despite the widespread adoption of open source software (OSS), its sustainability remains a critical concern, particularly in light of security vulnerabilities and the often inadequate end-of-service (EoS) processes for OSS projects as they decline. Existing models of OSS community participation, like the Onion model and the episodic contribution model, offer valuable insights but are fundamentally incompatible and fail to provide a comprehensive picture of contributor engagement with OSS projects. This paper addresses these gaps by proposing the CROSS model, a novel contributor-project interaction lifecycle model for open source, which delineates the various lifecycle stages of contributor-project interaction along with the driving and retaining forces pertinent to each stage. By synthesizing existing research on OSS communities, organizational behavior, and human resource development, it explains a range of archetypal cases of contributor engagement and highlights research gaps, especially in EoS/offboarding scenarios. The CROSS model provides a foundation for understanding and enhancing the sustainability of OSS projects, offering a robust foundation for future research and practical application.
@@ 2.2 Synthesis of the CROSS Model
The interaction between OSS projects and contributors is inherently complex, with each interaction being unique. However, there are common patterns that can be used to form a conceptual model of project-contributor engagement. While the idea of an OSS contributor lifecycle exists implicitly and different phases of this lifecycle have been explored in various studies, including but not limited to the ones noted above, it has yet to be formalized into a comprehensive model that captures the full range of interactions. To develop this model, we began by identifying the distinct phases of interaction between contributors and OSS projects, drawing from existing research and theories that describe various forms of contributor engagement. We synthesized these insights to build a model that describes the complete lifecycle of a contributor’s engagement with OSS projects and communities - the CROSS model - with each phase informed by relevant theoretical frameworks.
@@ 2.3.1 Phases:
    Onboarding: In this phase, the contributor attempts to get their contribution accepted in the project and might transition into being a member. This phase starts when the contributor decides to contribute to the project and ends when they have successfully transitioned into the project. During onboarding, the contributor learns the project’s contribution guidelines, prepares a contribution (often in the form of an issue or pull request), engages with project maintainers for feedback, and refines their contribution as needed. They may also interact with project members who mentor them. The mutual distance between the contributor and project members decreases as familiarity grows.

### v2

@@ Abstract
Despite the widespread adoption of open source software (OSS), its sustainability remains a critical concern, particularly in light of security vulnerabilities and the often inadequate end-of-service (EoS) processes for OSS projects as they decline. Existing models of OSS community participation, like the Onion model and the episodic contribution model, offer valuable insights but are fundamentally incompatible and fail to provide a comprehensive picture of contributor engagement with OSS projects. This paper addresses these gaps by proposing the CROSS model, a novel contributor-project interaction lifecycle model for open source, which delineates the various lifecycle stages of contributor-project interaction, along with the driving and retaining forces pertinent to each stage. By synthesizing existing research on OSS communities, organizational behavior, and human resource development, it explains a range of archetypal cases of contributor engagement and highlights research gaps, especially in EoS/offboarding scenarios. The CROSS model provides a foundation for understanding and enhancing the sustainability of OSS projects, offering a robust foundation for future research and practical application.
@@ 2.2 Synthesis of the CROSS Model
The interaction between OSS projects and contributors is inherently complex, with each interaction being unique. However, there are common patterns that can be used to form a conceptual model of project-contributor engagement. While the idea of an OSS contributor lifecycle exists implicitly, and different phases of this lifecycle have been explored in various studies, including but not limited to the ones noted above, it has yet to be formalized into a comprehensive model that captures the full range of interactions. To develop such a model, we began by identifying the distinct phases of interaction between contributors and OSS projects, drawing from existing research and theories that describe various forms of contributor engagement. We synthesized these insights to build a model that describes the complete lifecycle of a contributor’s engagement with OSS projects and communities - the CROSS model - with each phase informed by relevant theoretical frameworks.
@@ 2.3.1 Phases:
    Onboarding: In this phase, the contributor attempts to get their contribution accepted in the project and might transition into becoming a member. This phase starts when the contributor decides to contribute to the project and ends when they have successfully transitioned into the project. During onboarding, the contributor learns the project’s contribution guidelines, prepares a contribution (often in the form of an issue or pull request), engages with project maintainers for feedback, and refines their contribution as needed. They may also interact with project members who mentor them. The mutual distance between the contributor and project members decreases as familiarity grows.
@@ 5 Conclusion
Acknowledgement
The work was supported, in part, by Science Foundation Ireland grant 13/RC/2094 P2 to Lero.